a car traveling at 3.0 m/s has a constant acceleration of 4.0 m/s^2 after 2.0 seconds, the velocity is?
Found 2 solutions by Fombitz, Kasandra24:
Answer by Fombitz(32388)
Answer by Kasandra24(4)

at=Vf-Vi
Vf= Vi + at
Vf= 3 + 4(2)
=3 + 8
Final Velocity is 11 m/s